OpenStudy (anonymous):

Set up the integral to compute the volume of the solid that lies above the cone z =root(x^2+y^2) and below sphere x^2+y^2+z^2=4

OpenStudy (anonymous):

z=r^2 r^2+z^2=4 z=sqrt(4-r^2) use rdrd(thet)dz \[\int\limits_{0}^{2\pi} \int\limits_{0}^{\sqrt{2}} \int\limits_{\sqrt{4-r^2}}^{r}dzrdrd \theta\]

OpenStudy (anonymous):

the r and sqrt 4-r^2 shood be reversed sorry

OpenStudy (anonymous):

Thank you very much man.

OpenStudy (anonymous):

it might be right or i can do sperical coordinates

OpenStudy (anonymous):

i think the cylindrical coordinates here work just fine. thanks